The 17-kg wheel has an eccentric mass which places the center of mass G a distance r¯ = 142 mm away from the geometric center O. A constant couple M = 11.8 N∙m is applied to the initially stationary wheel, which rolls without slipping along the horizontal surface and enters the curve of radius R = 1110 mm. Determine the normal force NC under the wheel just before it exits the curve at C. The wheel has a rolling radius r = 100 mm and a radius of gyration kO = 141 mm.
M Answer: Nc = i
Answer #1

- Page noir Answer! 1. The objective is to determine the normal force No. under the wheel just before it exits the cuore at cpage noi.2. length of the path Be in 9 . 71 BCE RO 0111 : 0.87 length of BC = 90 0.87 90 0.87 :010 = 087 0: 8.7 h=on ģion 1,1Page ио3 33-94 2 = 8:5 (о? «- (-чѕ4)) + •5 o-ozo)(u - 1454) 33-42 = o-oss u° - 12 мет) + (-то (8° — 1us •1)] : o-ogs w -2N = [m xgx cos us] I fz tfi N: 17x9.81 x cos 45 + 319.34 + 672-43 page no. u. N: 235.16. kn. N: 0.235 kn 2. The normal force